# Spectrum-Aware Debiasing: A Modern Inference Framework with Applications to Principal Components Regression

### v4

@@ Abstract
Debiasing is a fundamental concept in high-dimensional statistics. While degrees-of-freedom adjustment is the state-of-the-art debiasing technique in high-dimensional linear regression, it largely remains limited to independent, identically distributed samples and sub-Gaussian covariates. These limitations hinder its wider practical use. In this paper, we break this barrier and introduce Spectrum-Aware Debiasing—-a novel inference method that applies to challenging high-dimensional regression problems with structured row-column dependencies, heavy tails, asymmetric properties, and latent low-rank structures. Our method achieves debiasing through a rescaled gradient descent step, where the rescaling factor is derived from the spectral properties of the sample covariance matrix. This spectrum-based approach enables accurate debiasing in much broader contexts. We study the common modern regime where the number of features and samples scale proportionally. We establish asymptotic normality of our proposed estimator (suitably centered and scaled) under various convergence notions when the covariates are right-rotationally invariant. Such designs have garnered significant recent attention due to their crucial role in compressed sensing. Furthermore, we devise a consistent estimator for its asymptotic variance.
Our work has two notable by-products: first, Spectrum-Aware Debiasing rectifies the bias in principal components regression (PCR), providing the first debiased PCR estimator in high dimensions. Second, we introduce a principled test for checking the presence of alignment between the signal and the eigenvectors of the sample covariance matrix. This test is independently valuable for statistical methods developed using approximate message passing, leave-one-out, or convex Gaussian min-max theorems. We demonstrate the utility of our method through diverse simulated and real data experiments. On the technical front, we connect approximate message passing algorithms with debiasing and provide the first proof of the Cauchy property of vector approximate message passing (V-AMP).
62E20,
62F12,
High-dimensional inference,
Debiasing,
Spectral Properties,
Principal Components Regression,
Debiased PCR,
Right-rotationally invariant designs,
@@ 1 Introduction
$\hat{\bm{\beta}}^{u}=\hat{\bm{\beta}}+\frac{1}{\widehat{\mathsf{adj}}}\bm{M}\mathbf{X}^{\top}(\mathbf{y}-\mathbf{X}\hat{\bm{\beta}}),$ (2)
@@
Classical statistics textbooks tell us that when the dimension $p$ is fixed and the sample size $n$ approaches infinity, the debiased estimator $\hat{\bm{\beta}}^{u}$ reduces to the well-known one-step estimator. In this case, Gaussianity of $\hat{\bm{\beta}}^{u}-\bm{\beta}^{\star}$ follows from [94, Theorem 5.45] by choosing $\mathbf{M}=(\mathbf{X}^{\top}\mathbf{X})^{-1}$, the inverse of the sample covariance matrix, and $\widehat{\mathsf{adj}}=1$, requiring no adjustment. Early work on ultra high-dimensional problems ($p\gg n$) [93, 105, 55, 16] established that when the signal $\bm{\beta}^{\star}$ is sufficiently sparse, the Lasso can be debiased by taking $\mathbf{M}$ as suitable “high-dimensional” substitutes of $(\mathbf{X}^{\top}\mathbf{X})^{-1}$ and setting $\widehat{\mathsf{adj}}=1$. However, later work uncovered that an adjustment of $\widehat{\mathsf{adj}}<1$ is necessary to relax sparsity assumptions on $\bm{\beta}^{\star}$ or to debias general regularized estimators beyond the Lasso. For instance, [53, 11] established that when the signal is not sufficiently sparse, the adjustment for the Lasso should be $\widehat{\mathsf{adj}}=1-\hat{s}/n$ with $\mathbf{M}=\bm{\Sigma}^{-1}$, where $\hat{s}$ denotes the number of non-zero entries in $\hat{\bm{\beta}}$. This correction term was named the “degrees-of-freedom adjustment” since $\hat{s}$ corresponds to the degrees-of-freedom of the estimator $\hat{\bm{\beta}}$ [110].
@@
Figure 1: Histograms of empirical distribution of $(\hat{\tau}_{*}^{-1/2}(\hat{{\beta}}^{u}_{i}-{\beta}^{\star}_{i}))_{i=1}^{p}$ comparing Degrees-of-Freedom Debiasing [12] with our Spectrum-Aware Debiasing, where $\hat{\bm{\beta}}^{u}$ is the debiased Elastic-Net estimator with tuning parameters $\lambda_{1}=1,\lambda_{2}=0.1$. The first row uses the Degrees-of-Freedom Debiasing from [12] with $\mathbf{M}=\mathbf{I}_{p}$ (denoted DF). The second row uses our Spectrum-Aware Debiasing (denoted SA) as in Definition 3.4. Entries of the signal $\bm{\beta}^{\star}$ are i.i.d. draws from $0.24\cdot N(-20,1)+0.06\cdot N(10,1)+0.7\cdot\delta_{0}$ where $\delta_{0}$ is Dirac-delta function at 00. Thereafter, the signal is fixed and we generate the responses using $\mathbf{y}=\mathbf{X}\bm{\beta}^{\star}+\bm{\varepsilon}$ where $\epsilon_{i}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}N(0,1)$. The solid black curve indicates a normal density fitted to the blue histograms whereas the dotted black line indicates the empirical mean corresponding to the histogram. See the corresponding QQ plot in Figure 5 from Appendix. All designs are re-scaled so that the average of eigenvalues of $\mathbf{X}^{\top}\mathbf{X}$ is 1. The design matrices are of shape $n=500,p=1000$.
@@
Degrees-of-Freedom Debiasing introduced a novel perspective. However, it relied on some strict assumptions, namely independent and identically distributed (i.i.d.) data with Gaussian covariates that follow $\mathbf{X}_{i}\sim\mathcal{N}(\bm{0},\bm{\Sigma})$. Furthermore, Degrees-of-Freedom Debiasing used $\mathbf{M}=\bm{\Sigma}^{-1}$; thus, even when the i.i.d. assumption holds, implementing this estimator required knowledge of the true covariance matrix $\bm{\Sigma}$ or an accurate estimate. In summary, Degrees-of-Freedom Debiasing suffers three key limitations: (i) it is restricted to Gaussian type distributions (see [43] for extension to sub-Gaussians), thus failing to capture heavy-tailed or asymmetrically distributed covariates; (ii) it is ineffective in scenarios with heterogeneity or dependency among samples, and (iii) it faces challenges in choosing $\mathbf{M}$ when precise estimates of $\bm{\Sigma}$ are unavailable. These limitations restrict the broader applicability of Degrees-of-Freedom Debiasing to real-world settings that violate these strict assumptions.
@@
We exemplify this issue in Figure 1, where we consider the following design distributions: (i) $\mathsf{MatrixNormal}$: $\mathbf{X}$ drawn from a matrix normal distribution with row and column correlations, i.e. $\mathbf{X}\sim N(\rm{0},\mathbf{\Sigma}^{\mathrm{(col)}}\otimes\mathbf{\Sigma}^{\mathrm{(row)}}),$ where $\mathbf{\Sigma}^{\mathrm{(col)}}_{ij}=0.5^{|i-j|}$ and $\bm{\Sigma}^{\mathrm{(row)}}$ follows an inverse-Wishart distribution with identity scale and degrees-of-freedom $1.1p$, which is chosen to be close to $p$ to encourage heavy-tails in the covariates; (ii) $\mathsf{Spiked}$: $\mathbf{X}$ contains latent structure, i.e. $\mathbf{X}=\alpha\cdot\mathbf{V}\mathbf{W}^{\top}+n^{-1}N(\rm{0},\mathbf{I}_{n}\otimes\mathbf{I}_{p})$ where $\alpha=10$ and $\mathbf{V}\in\mathbb{R}^{n\times m},\mathbf{W}\in\mathbb{R}^{p\times m}$ are drawn randomly from Haar matrices of dimensions $n,p$, and then we retain $m=50$ columns; (iii) $\mathsf{LNN}$: $\mathbf{X}$ formed by product of multiple random matrices (see [44, 45] for connections to linear neural networks), i.e. $\mathbf{X}=\mathbf{X}_{1}\cdot\mathbf{X}_{2}\cdot\mathbf{X}_{3}\cdot\mathbf{X}_{4}$ where $\mathbf{X}_{i}$’s have i.i.d. entries from $N(\rm{0},1)$; (iv) $\mathsf{VAR}$: rows of $\mathbf{X}$ drawn from a vector time series with the $i$-th row given by $\mathbf{X}_{i,\bullet}=\sum_{k=1}^{\tau\vee i}\alpha_{k}\mathbf{X}_{i-k,\bullet}+\bm{\varepsilon}_{i}$ where $\tau=3,\alpha=\quantity(0.4,0.08,0.04)$ and $\bm{\varepsilon}_{i}\sim N(\mathbf{0},\mathbf{\Sigma})$ with $\mathbf{\Sigma}$ drawn from an inverse-Wishart distribution with the same parameters as in (i); (v) $\mathsf{Mult}$-$\mathsf{t}$: rows of $\mathbf{X}$ drawn independently from a multivariate t-distribution with identity scale and degrees-of-freedom $3$. Figure 1 plot histograms of the empirical distribution of $\hat{\bm{\beta}}^{u}-\bm{\beta}^{\star}$ scaled by an estimate $\hat{\tau}_{*}$ of its standard deviation. The topmost panel uses the Degrees-of-Freedom Debiasing formula for $\hat{\bm{\beta}}^{u}$. We observe that the histograms in this panel deviate substantially from the overlaid standard Gaussian density. Degrees-of-Freedom Debiasing thus fails in these challenging settings. To underscore the difficulties posed by these examples, note that cases (i)-(iv) involve non-i.i.d. designs and (i),(iv),(v) involve heavy-tailed covariates.
@@
In this paper, we propose a new debiasing formula that addresses the shortcomings of previous techniques and enables accurate debiasing in the aforementioned settings. To develop our method, we leverage the insight that a debiasing procedure effective for a wide range of scenarios must thoughtfully utilize the spectral characteristics of the data. To accomplish this, we explore an alternative path for modeling the randomness in the design. Instead of assuming that the rows of the design are i.i.d. Gaussian vectors, we require that the singular value decomposition of $\mathbf{X}$ satisfies certain natural structure that allows dependence among samples and potentially heavy-tailed distributions. Specifically, we assume that $\mathbf{X}$ is right-rotationally invariant (formally defined in Definition 2.1).
This class of designs has been widely studied in signal processing, information theory, and statistical physics [85, 64, 74, 86, 26, 31, 40, 87, 88, 95, 41, 62, 59, 97]. They serve as important prototypes for understanding fundamental high-dimensional phenomena in compressed sensing. Furthermore, recent advancements indicate that a wide variety of covariate distributions fall within the same universality class as these designs, provided that the eigenvectors of the sample covariance matrix are adequately “generic” [23, 27, 96]. Thus, the range of distributions covered by these designs is much larger than those considered in prior debiasing literature. In particular, this includes designs (i)-(v) from Figure 1.
@@
$\hat{\bm{\beta}}^{u}=\hat{\bm{\beta}}+\widehat{\mathsf{adj}}^{-1}\mathbf{X}^{\top}(\mathbf{y}-\mathbf{X}\hat{\bm{\beta}}),$ (3)
@@
Here, $\{d_{i}^{2}\}_{1\leq i\leq p}$ represents the eigenvalues of the sample covariance matrix $\mathbf{X}^{\top}\mathbf{X}$, and $h^{\prime\prime}$ denotes the second derivative of the penalty function $h$ used in calculating the regularized estimator $\hat{\bm{\beta}}^{u}$. At points of non-differentiability, we extend $h^{\prime\prime}$ by +$\infty$+ ∞ (cf. Section 2.2). The solution $\widehat{\mathsf{adj}}$ of (4) is unique for any $p\geq 1$ under mild assumptions (cf. Section 3.1). We refer to $\widehat{\mathsf{adj}}$ as “Spectrum-Aware adjustment” and the debiasing approach in (3) as “Spectrum-Aware Debiasing” since $\widehat{\mathsf{adj}}$ depends on the eigenvalues of $\mathbf{X}^{\top}\mathbf{X}$. Figure 1 illustrates the efficacy of our method. The second panel shows the empirical distribution of $\hat{\bm{\beta}}^{u}-\bm{\beta}^{\star}$, scaled by an appropriate estimate of its standard deviation, when $\hat{\bm{\beta}}^{u}$ is given by our Spectrum-Aware formula (3). Note the remarkable agreement with the overlaid standard Gaussian density. We should mention that, unlike Degrees-of-Freedom Debiasing, we do not need an estimate of $\bm{\Sigma}$, yet we can handle various correlation structures among the features, as demonstrated in Figure 1.
@@
    We establish that our proposed debiasing formula is well-defined, that is, (4) admits a unique solution (Proposition 3.1). Then we establish that $\hat{\bm{\beta}}^{u}-\bm{\beta}^{\star}$, with this choice of $\widehat{\mathsf{adj}}$, converges to a mean-zero Gaussian with some variance $\tau_{\star}$ in a Wasserstein-2 sense (Theorem 3.1; Wasserstein-2 convergence notion introduced in Definition 2.2). Under an exchangeability assumption on $\bm{\beta}^{\star}$, we strengthen this result to convergence guarantees on finite-dimensional marginals of $\hat{\bm{\beta}}^{u}-\bm{\beta}^{\star}$ (Corollary 3.8).
@@
    We develop a consistent estimator for $\tau_{*}$ (Theorem 3.1) by developing new algorithmic insights and new proof techniques that can be of independent interest in the context of vector approximate message passing algorithms [74, 77, 37] (details in Section 5.2).
@@
    Theoretically, we establish rigorously the risk of regularized estimators under right-rotationally invariant designs, a class much larger than Gaussian designs (Theorem 5.1). One should compare this theorem to [7] that developed a risk characterization of the Lasso in high dimensions under Gaussian design matrices.
@@
In the remaining Introduction, we walk the readers through some important discussion points, before we delve into our main results. In Section 1.1, we provide some intuition for our Spectrum-Aware construction using the example of the ridge estimator, since it admits a closed form and is simple to study. In Section 1.2, we discuss right-rotationally invariant designs and related literature, with the goal of motivating why this assumption is natural for debiasing, and emphasizing its degree of generality compared to prior Gaussian assumptions. In Section 1.3, we describe how the debiasing methods tend to fail when the design $\mathbf{X}$ contains outlier eigenvalues and/or the signal aligns with some eigenvectors of $\mathbf{X}$. In Section 1.4, we discuss a novel PCR-Spectrum-Aware Debiasing approach which addresses the aforementioned two issues and an associated hypothesis test for alignment between signal and principal components.

### v6

@@ Abstract
Debiasing is a fundamental concept in high-dimensional statistics. While degrees-of-freedom adjustment is the state-of-the-art debiasing technique in high-dimensional linear regression, it largely remains limited to independent, identically distributed samples and sub-Gaussian covariates. These limitations hinder its wider practical use. In this paper, we break this barrier and introduce Spectrum-Aware Debiasing—-a novel inference method that applies to challenging high-dimensional regression problems with structured row-column dependencies, heavy tails, asymmetric properties, and latent low-rank structures. Our method achieves debiasing through a rescaled gradient descent step, where the rescaling factor is derived from the spectral properties of the sample covariance matrix. This spectrum-based approach enables accurate debiasing in much broader contexts. We study the common modern regime where the number of features and samples scale proportionally. We establish asymptotic normality of our proposed estimator (suitably centered and scaled) under various convergence notions when the covariates are right-rotationally invariant. We further prove a spectral universality result, extending our guarantees to a much broader class of covariate distributions. Furthermore, we devise a consistent estimator for the asymptotic variance.
Our work has two notable by-products: first, Spectrum-Aware Debiasing rectifies the bias in principal components regression (PCR), providing the first debiased PCR estimator in high dimensions. Second, we introduce a principled test for checking the presence of alignment between the signal and the eigenvectors of the sample covariance matrix. This test is independently valuable for statistical methods developed using approximate message passing, leave-one-out, random matrix theory, or convex Gaussian min-max theorems. We demonstrate the utility of our method through diverse simulated and real data experiments.
@@ 1 Introduction
Classical statistics textbooks tell us that when the dimension $p$ is fixed and the sample size $n$ approaches infinity, the debiased estimator $\hat{\bm{\beta}}^{u}$ reduces to the well-known one-step estimator. In this case, Gaussianity of $\hat{\bm{\beta}}^{u}-\bm{\beta}^{\star}$ follows from [110, Theorem 5.45] by choosing $\mathbf{M}=(\mathbf{X}^{\top}\mathbf{X})^{-1}$, the inverse of the sample covariance matrix, and $\widehat{\mathsf{adj}}=1$, requiring no adjustment. Early work on ultra high-dimensional problems ($p\gg n$) [109, 123, 63, 18] established that when the signal $\bm{\beta}^{\star}$ is sufficiently sparse, the Lasso can be debiased by taking $\mathbf{M}$ as suitable “high-dimensional” substitutes of $(\mathbf{X}^{\top}\mathbf{X})^{-1}$ and setting $\widehat{\mathsf{adj}}=1$. However, later work uncovered that an adjustment of $\widehat{\mathsf{adj}}<1$ is necessary to relax sparsity assumptions on $\bm{\beta}^{\star}$ or to debias general regularized estimators beyond the Lasso. For instance, [61, 12] established under the proportional regime ($n/p\to w>0$) that when the signal is not sufficiently sparse, the adjustment for the Lasso should be $\widehat{\mathsf{adj}}=1-\hat{s}/n$ with $\mathbf{M}=\bm{\Sigma}^{-1}$, where $\hat{s}$ denotes the number of non-zero entries in $\hat{\bm{\beta}}$ and $\bm{\Sigma}$ is the covariance matrix of i.i.d. Gaussian rows of $\mathbf{X}$. This correction term was named the “degrees-of-freedom adjustment” since $\hat{s}$ corresponds to the degrees-of-freedom of the estimator $\hat{\bm{\beta}}$ [126].
@@
Figure 1: Histograms of empirical distribution of $(\hat{\tau}_{*}^{-1/2}(\hat{{\beta}}^{u}_{i}-{\beta}^{\star}_{i}))_{i=1}^{p}$ comparing Degrees-of-Freedom Debiasing [13] with our Spectrum-Aware Debiasing, where $\hat{\bm{\beta}}^{u}$ is the debiased Elastic-Net estimator with tuning parameters $\lambda_{1}=1,\lambda_{2}=0.1$. The first row uses the Degrees-of-Freedom Debiasing from [13] with $\mathbf{M}=\mathbf{I}_{p}$ (denoted DF). The second row uses our Spectrum-Aware Debiasing (denoted SA) as in Definition 3.5. Entries of the signal $\bm{\beta}^{\star}$ are i.i.d. draws from $0.24\cdot N(-20,1)+0.06\cdot N(10,1)+0.7\cdot\delta_{0}$ where $\delta_{0}$ is Dirac-delta function at 0. Thereafter, the signal is fixed and we generate the responses using $\mathbf{y}=\mathbf{X}\bm{\beta}^{\star}+\bm{\varepsilon}$ where $\varepsilon_{i}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}N(0,1)$. The solid black curve indicates a normal density fitted to the blue histograms whereas the dotted black line indicates the empirical mean corresponding to the histogram. See the corresponding $Q$ plot in Figure 5 from Appendix. All designs are re-scaled so that the average of eigenvalues of $\mathbf{X}^{\top}\mathbf{X}$ is 1. The design matrices are of shape $n=500,p=1000$.
@@
Degrees-of-Freedom Debiasing introduced a novel perspective. However, it relied on some strict assumptions, namely independent and identically distributed (i.i.d.) data with Gaussian covariates that follow $\mathbf{X}_{i}\sim N(\bm{0},\bm{\Sigma})$. Furthermore, Degrees-of-Freedom Debiasing used $\mathbf{M}=\bm{\Sigma}^{-1}$; thus, even when the i.i.d. assumption holds, implementing this estimator required knowledge of the true covariance matrix $\bm{\Sigma}$ or an accurate estimate. In summary, Degrees-of-Freedom Debiasing suffers three key limitations: (i) it is restricted to Gaussian type distributions (see [53] for extension to sub-Gaussians), thus failing to capture heavy-tailed or asymmetrically distributed covariates; (ii) it is ineffective in scenarios with heterogeneity or dependency among samples, and (iii) it faces challenges in choosing $\mathbf{M}$ when precise estimates of $\bm{\Sigma}$ are unavailable. These limitations restrict the broader applicability of Degrees-of-Freedom Debiasing to real-world settings that violate these strict assumptions.
@@
We exemplify this issue in Figure 1, where we consider the following design distributions: (i) $\mathsf{MatrixNormal}$: $\mathbf{X}$ drawn from a matrix normal distribution with row and column correlations, i.e. $\mathbf{X}\sim N(\rm{0},\mathbf{\Sigma}^{\mathrm{(col)}}\otimes\mathbf{\Sigma}^{\mathrm{(row)}}),$ where $\mathbf{\Sigma}^{\mathrm{(col)}}_{ij}=0.5^{|i-j|}$ and $\bm{\Sigma}^{\mathrm{(row)}}$ follows an inverse-Wishart distribution with identity scale and degrees-of-freedom $1.1p$, which is chosen to be close to $p$ to encourage heavy-tails in the covariates; (ii) $\mathsf{Spiked}$: $\mathbf{X}$ contains latent structure, i.e. $\mathbf{X}=\alpha\cdot\mathbf{V}\mathbf{W}^{\top}+n^{-1}N(\rm{0},\mathbf{I}_{n}\otimes\mathbf{I}_{p})$ where $\alpha=10$ and $\mathbf{V}\in\mathbb{R}^{n\times m},\mathbf{W}\in\mathbb{R}^{p\times m}$ are drawn randomly from Haar matrices of dimensions $n,p$, and then we retain $m=50$ columns; (iii) $\mathsf{LNN}$: $\mathbf{X}$ formed by product of multiple random matrices (see [54] for connections to linear neural networks), i.e. $\mathbf{X}=\mathbf{X}_{1}\cdot\mathbf{X}_{2}\cdot\mathbf{X}_{3}\cdot\mathbf{X}_{4}$ where $\mathbf{X}_{i}$’s have i.i.d. entries from $N(\rm{0},1)$; (iv) $\mathsf{VAR}$: rows of $\mathbf{X}$ drawn from a vector time series with the $i$-th row given by $\mathbf{X}_{i,\bullet}=\sum_{k=1}^{\tau\vee i}\alpha_{k}\mathbf{X}_{i-k,\bullet}+\bm{\varepsilon}_{i}$ where $\tau=3,\alpha=\quantity(0.4,0.08,0.04)$ and $\bm{\varepsilon}_{i}\sim N(\mathbf{0},\mathbf{\Sigma})$ with $\mathbf{\Sigma}$ drawn from an inverse-Wishart distribution with the same parameters as in (i); (v) $\mathsf{Mult}$-$\mathsf{t}$: rows of $\mathbf{X}$ drawn independently from a multivariate t-distribution with identity scale and degrees-of-freedom $3$. Figure 1 plot histograms of the empirical distribution of $\hat{\bm{\beta}}^{u}-\bm{\beta}^{\star}$ scaled by an estimate $\hat{\tau}_{*}$ of its standard deviation. The topmost panel uses the Degrees-of-Freedom Debiasing formula for $\hat{\bm{\beta}}^{u}$. We observe that the histograms in this panel deviate substantially from the overlaid standard Gaussian density. Degrees-of-Freedom Debiasing thus fails in these challenging settings.
To underscore the difficulties posed by these examples, note that cases (i)-(iv) involve non-i.i.d. designs and (i),(iv),(v) involve heavy-tailed covariates. As discussed later, the failure observed is primarily attributable to these structural deviations rather than finite-sample effects.
@@
In this paper, we propose a new debiasing formula that addresses the shortcomings of previous techniques and enables accurate debiasing in the aforementioned settings. To develop our method, we leverage the insight that a debiasing procedure effective for a wide range of scenarios must thoughtfully utilize the spectral characteristics of the data. To accomplish this, we explore an alternative path for modeling the randomness in the design. Instead of assuming that the rows of the design are i.i.d. Gaussian vectors, we require that the singular value decomposition of $\mathbf{X}$ satisfies certain natural structure that allows dependence among samples and potentially heavy-tailed distributions. Specifically, we assume that $\mathbf{X}$ is right-rotationally invariant (Definition 2.1).
Right-rotationally invariant designs have been widely studied in signal processing, information theory, statistical physics, and high-dimensional statistics [100, 78, 88, 101, 32, 38, 48, 102, 103, 111, 49, 74, 69, 113, 77], and they serve as useful prototypes for fundamental high-dimensional phenomena in compressed sensing. Roughly speaking, if the right singular vectors $\mathbf{O}$ of a design matrix $\mathbf{X}$ are Haar-distributed, then $\mathbf{X}$ lies in the class of right-rotationally invariant designs, regardless of distribution of its eigenvalues. This generality lets us handle design distributions not covered by degree-of-freedom–based debiasing methods, including designs (i)–(v) in Figure 1. Since the right-rotational invariance assumption preserves the spectral information of $\mathbf{X}^{\top}\mathbf{X}$, we also expect methods developed under this assumption to exhibit improved robustness when applied to real-data designs. This is illustrated in Figure 3, where we evaluate our PCR–Spectrum–Aware Debiasing method based on right-rotational invariance assumption on six real datasets spanning image data, financial data, socio-economic data and so forth. Furthermore, recent advances indicate that a wide variety of covariate distributions fall within the same universality class as right-rotationally invariant designs, provided the eigenvectors of the sample covariance are sufficiently “generic,” even if not exactly Haar [29, 34, 112]. In Appendix, Section F.1, we extend our results to this broader “spectral universality class” [34, 112], which encompasses right-rotationally invariant designs, i.i.d. designs and their linear transforms as well as other challenging design distributions. A more detailed discussion of the technical challenges associated with right-rotational invariance, along with a review of relevant prior work, is provided in Appendix, Appendix E.
@@
[…]
@@
Here, $\{d_{i}^{2}\}_{1\leq i\leq p}$ represents the eigenvalues of the sample covariance matrix $\mathbf{X}^{\top}\mathbf{X}$, and $h^{\prime\prime}$ denotes the second derivative of the penalty function $h$ used in calculating the regularized estimator $\hat{\bm{\beta}}^{u}$. At points of non-differentiability (e.g. $x=0$ for Lasso and Elastic Net), we extend $h^{\prime\prime}$ by $\infty$ (cf. Section 2.2). The solution $\widehat{\mathsf{adj}}$ of (4) is unique for any $p\geq 1$ under mild assumptions (cf. Section 3.1). We refer to $\widehat{\mathsf{adj}}$ as the “Spectrum-Aware adjustment” and the debiasing approach in (3) as “Spectrum-Aware Debiasing” since $\widehat{\mathsf{adj}}$ depends on the eigenvalues of $\mathbf{X}^{\top}\mathbf{X}$. Figure 1 illustrates the efficacy of our method. The second panel shows the empirical distribution of $\hat{\bm{\beta}}^{u}-\bm{\beta}^{\star}$, scaled by an appropriate estimate of its standard deviation, when $\hat{\bm{\beta}}^{u}$ is given by our Spectrum-Aware formula (3). Note the remarkable agreement with the overlaid standard Gaussian density. We emphasize that this debiasing formula relies on a fundamentally different structured dependency assumption than Degrees-of-Freedom Debiasing. The type of dependency it captures is incomparable to that of anisotropic Gaussians.222for the analogue of anisotropic Gaussian-type dependence in the context of right rotationally invariant designs, see Appendix I. Nevertheless, our method can capture quite diverse dependency structures, as demonstrated in Figure 1. Crucially, it operates without requiring an estimate of the population feature covariance matrix.
@@
    We establish that our proposed debiasing formula is well-defined, that is, (4) admits a unique solution (Proposition 3.1). Then we establish that $\hat{\bm{\beta}}^{u}-\bm{\beta}^{\star}$, with this choice of $\widehat{\mathsf{adj}}$, converges to a mean-zero Gaussian with some variance $\tau_{*}$ in a Wasserstein-2 sense (Theorem 3.1; Wasserstein-2 convergence notion introduced in Definition 2.2). Under an exchangeability assumption on $\bm{\beta}^{\star}$, we strengthen this result to convergence guarantees on finite-dimensional marginals of $\hat{\bm{\beta}}^{u}-\bm{\beta}^{\star}$ (Corollary 3.10).
@@
    We develop a consistent estimator for $\tau_{*}$ (Theorem 3.1) by developing new algorithmic insights and new proof techniques that can be of independent interest in the context of vector approximate message passing algorithms [88, 92, 44] (details in Section C.2).
@@
    On the technical front, we rigorously characterize the risk of regularized estimators under right-rotationally invariant designs (cf. Theorem C.1), and extend these results to a broader spectral universality class (cf. Corollary G.3). We prove existence and uniqueness of the solution associated with our fixed-point equations under appropriate conditions (cf. Proposition 2.11). We establish the Cauchy convergence of VAMP iterates (cf. Corollary C.9). We further extend our results to the challenging case of the Lasso under suitable sparsity conditions (cf. Theorem F.3)—this requires substantial arguments beyond those for strongly convex penalties (cf. Appendix H). We note that analogs of the leave-one-out approach [80, 104, 11, 35, 98, 97, 24, 64] and Stein’s method [94, 23, 13, 4], both of which form fundamental proof techniques for Gaussian designs, are nonexistent or under-developed for rotationally invariant designs. Therefore, our approach adopts an algorithmic proof strategy inspired by prior work from the senior authors and others in the Gaussian case.
@@
In the remaining Introduction, we walk the readers through some important discussion points, before we delve into our main results. In Section 1.1, we provide some intuition for our Spectrum-Aware construction using the example of the ridge estimator, since it admits a closed form and is simple to study. In Section 1.2, we describe how the debiasing methods tend to fail when the design $\mathbf{X}$ contains outlier eigenvalues and/or the signal aligns with some eigenvectors of $\mathbf{X}$. In Section 1.3, we discuss a novel PCR-Spectrum-Aware Debiasing approach which addresses the aforementioned two issues and an associated hypothesis test for alignment between signal and principal components.